why are iron compounds more readily oxidised by air in an alkaline medium than in an acidic medium
1
Expert's answer
2012-09-14T09:26:33-0400
Iron compounds more readily oxidised by air in acidic medium,cause

Fe2+
+1/2 O2 + H2O + 1ê =Fe3+ 2 OH-

In alkaline medium equilibrium reaction
will shift toward the formation of primary products,so better is acidic
medium

FeSO4+O2+H2SO4=Fe2(SO4)3+H2O